When production function of a firm is Q = 20K^0.5L^0.5 , price of capital is 5 per unit and price of labour is 4 per unit . What will be the expansion path ratio for the firm?
1
Expert's answer
2021-09-19T18:34:31-0400

Given

"Q = 20 K^{0.5}L^{0.5}"

"MPL = \frac{dq}{dl}\\

= 20\times 0.5×L^{-0.5} ×K^{0.5}\\

=10×L^{-0.5} ×K^{0.5}"


"MPK = \frac{dq}{dl}\\

= 20\times 0.5×L^{0.5} ×K^{-0.5}\\

=10×L^{0.5} ×K^{-0.5}"


Now expansion path = MPL/Price of labor= MPK/Price of capital

"\frac{[10×L^{-0.5} ×K^{0.5}]}{ 4} =\frac{ [ 10×L^{0.5}×K^{-0.5} ]}{5}\\

\frac{K}{4} =\frac{L}{5}\\

K=\frac{4L}{5}"